OpenStudy (anonymous):

Find the inverse function of f(t)=75.457-0.707t

OpenStudy (aravindg):

what do u think it will be

OpenStudy (aravindg):

?

OpenStudy (aravindg):

helo?

OpenStudy (auctoratrox):

do you know how to do this? when you get the answer, it's f^(-1)(t) = (75.457-t)/0.707. This graph is a reflection of the original function over the line y = x.

OpenStudy (australopithecus):

To find the inverse function switch y with t and t with y and solve for y so y = 75.457-0.707t becomes t = 75.457-0.707y = 0.707y = 75.457 - t = y = (75.457 - t)/0.707 thus f^(-1)(t) = (75.457 - t)/0.707

OpenStudy (australopithecus):

I explain inverse functions in this post if you are interested http://openstudy.com/users/australopithecus#/updates/4fae8ea0e4b059b524f9c5e3